Find // Vivi äysin (x2 + y²)dxdy
Answer #1

Solution 1/3 0 substitute u= 2 7 du = 22 - 27 3 y?+ 1/35 y 12 [ ay sin (a +92) dalay x² + y2 when a=0, u=y XA , U= 9²4 35 Yh/T s cas (4) (²) du Al (siniusban 2 2 (sinca) - Sin ()) = 1 - 3 (2713 y cos 142) dy 0 substitute a u=y? du =24 a dy when y=0,